# Removing Undesirable Concepts in Text-to-Image Generative Models with Learnable Prompts

### v1

@@ 4.2 Knowledge Transfer and Removal with Prompt - KPOP
$\underset{\theta^{{}^{\prime}}:\|\theta^{{}^{\prime}}-{\theta}_{k}^{{}^{\prime}}\|_{2}\leq\rho}{\min}\;\mathbb{E}_{c_{e}\in\mathbf{E}}\left[\underbrace{\left\|\epsilon_{\theta^{{}^{\prime}}}(c_{e})-\epsilon_{\theta}(c_{n})\right\|_{2}^{2}}_{L1}+\lambda\underbrace{\left\|\epsilon_{\theta^{{}^{\prime}}}(c_{e},\mathbf{p}_{k+1})-\epsilon_{\theta}(c_{e})\right\|_{2}^{2}}_{L2}\right],$ (7)
where we again use one-step gradient descent to update $\theta^{{}^{\prime}}$.
$\theta^{\prime}_{k+1}=\theta^{{}^{\prime}}_{k}-\eta\nabla_{\theta^{{}^{\prime}}}\mathcal{L}_{r}\left(\theta^{{}^{\prime}}\right),$
@@
with $\mathcal{L}_{r}\left(\theta^{{}^{\prime}}\right)=\mathbb{E}_{c_{e}\in\mathbf{E}}\left[\left\|\epsilon_{\theta^{{}^{\prime}}}(c_{e})-\epsilon_{\theta}(c_{n})\right\|_{2}^{2}+\lambda\left\|\epsilon_{\theta^{{}^{\prime}}}(c_{e},\mathbf{p}_{k+1})-\epsilon_{\theta}(c_{n})\right\|_{2}^{2}\right]$.
Minimizing the loss $L_{1}$ in Eq. 7 directly eliminates the undesirable concepts associated with $c_{e}$ by mapping it to a neutral or null concept $c_{n}$. However, it can be seen that optimizing solely for $L_{1}$ does not consider the impact on other concepts. Consequently, there might be many solutions $\theta^{{}^{\prime}}$ of $L_{1}$ that may successfully remove the undesirable concepts, yet they could inadvertently compromise other concepts due to parameter space overlap among concepts.
Minimizing the loss $L_{2}$ in Eq. 7 provides two benefits. First, it serves as a crucial regularization component, narrowing down the solution space of $L_{1}$. Specifically, as a result of the knowledge transfer stage, $\epsilon_{\theta^{{}^{\prime}}_{k}}(c_{e},\mathbf{p}_{k+1})\approx\epsilon_{\theta}(c_{e})$ with the knowledge of undesirable concepts attributed to the prompt $\mathbf{p}_{k+1}$. Thus, minimizing $L_{2}$ in Eq. 7 narrows down the solution space of $L_{1}$ to those that not only map the erased concepts to the neutral concept but also preserve the knowledge of the prompt. Notably, compared to the prior work (Orgad et al., 2023), which employs L2 regularization directly to the parameter space $\left\|\theta^{{}^{\prime}}-\theta\right\|_{2}^{2}$, our method incorporates regularization on the output space. This implicit consideration takes into account the varying importance of different parameters to the output, providing a more nuanced approach.
Second, minimizing the loss $L_{2}$ also helps transfer the dependency of generating undesirable concepts from the textual input $c_{e}$ to the prompt $\mathbf{p}$. The model cannot generate the undesirable concepts without the presence of the prompt $\mathbf{p}$, i.e., $\epsilon_{\theta^{{}^{\prime}}}(c_{e})\neq\epsilon_{\theta}(c_{e})$ while $\epsilon_{\theta^{{}^{\prime}}}(c_{e},\mathbf{p})\approx\epsilon_{\theta}(c_{e})$. As demonstrated in Section 5.5, we show that the knowledge of undesirable concepts is gradually transferred to the prompt $\mathbf{p}$ as expected.
4.3 Cross-Attention with Prompt
In this section, we will delve into discussing two different mechanisms for injecting the prompt into the cross-attention module, known as the concatenative mechanism and the additive mechanism. Their basic operations can be found in Table 1. In text-to-image diffusion models, the cross-attention layers are positioned to integrate textual embeddings into visual generation to regulate the output’s concept. Hence, these specific layers are as the most suitable for prompt injection, aligning with our goal.
Concatenative Mechanism.
In this mechanism, the prompt is concatenated with the textual embedding $\mathbf{C}$ before being used as the key and value matrix inputs to the cross-attention module. Let $\mathbf{p}\in\mathbb{R}^{[1,m_{p},d_{p}]}$ be the additional learnable prompt, where $m_{p}=km_{c}$ is the prompt size and $d_{p}=d_{c}$ is the dimension of the prompt. The main difference compared to the original mechanism is the projected matrices $K$ and $V$ as shown in Table 1.
Softmax normalization is applied on the last dimension of the attention score matrix $\mathbf{A}$. In addition, the scaling factor $\sqrt{d}$ is used to prevent the attention score from being too small when the dimension of the latent vector is large (Vaswani et al., 2017).
One of the advantages of this mechanism is that adding the prompt does not change the mechanism of the cross-attention module, which means that there is no need for a new architecture and it does not interfere with the model’s ability to generate good content. We can use the same pre-trained model and only need to modify the corresponding cross-attention module in the codebase. With $\mathbf{p}=c$, the model can generate the same output as the original model.
Secondly, this mechanism allows us to utilize a larger prompt (theoretically, the prompt can be arbitrarily large). By using a larger prompt, we can either remove more undesirable content or preserve more desirable content. Experiments in Section 5.6.2 demonstrate that the performance of this mechanism is scalable with the size of the prompt.
Figure 1: Illustration of Cross Attention with additional prompt.
Limitation of the concatenative mechanism: The main limitation of the concatenative mechanism is that it relies on softmax normalization to distribute the attribution from the additional prompt to the entire textual path. This issue is also illustrated in Figure 1. Because of the linearity of the projection operation, the matrices $\mathbf{K}$ and $\mathbf{V}$ can be decomposed into two parts separately, the projections of original textual embedding $\mathbf{C}$ (blue color) and prompt $\mathbf{P}$ (orange color), i.e., $\mathbf{K}=[\mathbf{W}_{k}\mathbf{C},\mathbf{W}_{k}\mathbf{P}],\mathbf{V}=[\mathbf{W}_{v}\mathbf{C},\mathbf{W}_{v}\mathbf{P}]$. As a result, the dot-product between the query $\mathbf{Q}$ and the key $\mathbf{K}$ can also be decomposed into two disjointed parts, i.e., $\mathbf{Q}\mathbf{K}^{T}=[\mathbf{Q}(\mathbf{W}_{k}\mathbf{C})^{T},\mathbf{Q}(\mathbf{W}_{k}\mathbf{P})^{T}]$. Without the softmax normalization, the output of the cross-attention module will be just the sum of the two disjointed parts, which is not desirable. The softmax normalization applied on the last dimension of the dot-product score matrix $\mathbf{Q}\mathbf{K}^{T}$ helps to distribute the attribution from the prompt to the entire textual path, enabling the model to attend to the additional prompt better. However, as the prompt size increases, the softmax normalization will distribute the attribution from the prompt to the entire textual path more evenly, which can lead to the model’s inability to attend to the prompt effectively, as shown in Section 5.6.2.
Additive Mechanism.
This mechanism injects the additional prompt by directly adding it to the textual embedding $\mathbf{C}$ before being used as the key and value matrix inputs to the cross-attention module. This retains the same advantages as the concatenative mechanism, i.e., it does not change the mechanism of the cross-attention module. It also permits a deeper integration of the prompt into the textual path, which allows the model to attend to the prompt more effectively. However, a limitation of this mechanism is that it is not scalable since its size is fixed to the size of the textual embedding. We compare the performance of two mechanisms in Section 5.6.1.
Alternative Prompting Mechanism.
Beyond the two aforementioned mechanisms, we acknowledge that there are several potential prompting mechanisms that can be used to modify the cross-attention module. For example, we can inject the prompt before the text encoder, by using a learnable word embedding vector associated with a special token $S^{*}$ to represent the prompt as in Textual Inversion (Gal et al., 2022). We can also amortize the prompt by using a learnable function to generate the prompt from the textual embedding, i.e., $\mathbf{p}=f(\mathbf{c})$. We leave the exploration of these mechanisms for future work.
@@ 5.1 General Settings
Our experiments are conducted using Stable Diffusion (SD) version 1.4 as the foundation model. We employ the same setting across all methods, i.e., fine-tuning the model for 1000 steps with a batch size of 1, with the Adam optimizer with a learning rate of $1\mathrm{e}{-5}$. We benchmark our method against four baseline approaches, namely, the original pre-trained SD model, ESD (Gandikota et al., 2023), UCE (Gandikota et al., 2024), and Concept Ablation (CA) (Kumari et al., 2023). Our method (KPOP), in its default configuration, employs a prompt size of $k=10$ and utilizes the concatenative mechanism. The effectiveness of these specific settings is discussed in detail in the ablation study section and the supplementary material.
@@ 5.2 Erasing Object-Related Concepts
In this experiment, we investigate the ability of our method to erase object-related concepts from the foundation model, for example, erasing entire object classes such as ‘Cassette Player’ from the model. We choose Imagenette  [^1] https://github.com/fastai/imagenette which is a subset of the ImageNet dataset (Deng et al., 2009) which comprises 10 easily recognizable classes, including ‘Cassette Player’, ‘Chain Saw’, ‘Church’, ‘Gas Pump’, ‘Tench’, ‘Garbage Truck’, ‘English Springer’, ‘Golf Ball’, ‘Parachute’, and ‘French Horn’.
Since the erasing performance when erasing a single class has been the main focus of previous work (Gandikota et al., 2023), we choose a more challenging setting where we erase a set of 5 classes simultaneously. Specifically, we generate 500 images for each class and employ the pre-trained ResNet-50 (He et al., 2016) to detect the presence of an object in the generated images. We use the two following metrics to evaluate the erasing performance: Erasing Success Rate (ESR-k): The percentage of all the generated images with "to-be-erased" classes where the object is not detected in the top-k predictions. Presevering Success Rate (PSR-k): The percentage of all the generated images with all other classes (i.e., "to-be-preserved") where the object is detected in the top-k predictions. This dual-metric evaluation provides a comprehensive assessment of our method’s ability to effectively erase targeted object-related concepts while also preserving relevant elements.
Quantitative Results.
@@
We select four distinct sets of 5 classes from the Imagenette set for erasure and present the outcomes in Table 2. First, we note that the average PSR-1 and PSR-5 scores across the four settings of the original SD model stand at 78.0% and 97.6%, respectively. These scores indicate that 78.0% of the generated images contain the object-related concepts which are subsequently detected in the top-1 prediction, and when checking the conepts in any of the top-5 predictions, this number increases to 97.6%. This underscores the original SD models’s ability to generate images with the anticipated object-related concepts.
@@
In term of erasing performance, it can be observed that all baselines achieve very high ESR-1 and ESR-5 scores, with the lowest ESR-1 and ESR-5 scores being 95.5% and 88.9% respectively. This indicates the effectiveness of these methods to erase object-related concepts, as only a very small proportion of the generated images contain the object-related concepts under subsequent detection. Notably, the UCE method can achieve 100% ESR-1 and ESR-5, which is the highest among the baselines. Our method achieves 99.2% ESR-1 and 97.3% ESR-5, which is much higher than the two baselines ESD and CA, and only slightly lower than the UCE method, which is designed specifically for erasing object-related concepts.
@@
However, despite the high erasing performance, the baselines suffer from a significant drop in preserving performance, with the lowest PSR-1 and PSR-5 scores being 41.2% and 56.1%, respectively. This suggests that the preservation task poses greater challenges than the erasing task, and the baselines are ineffective in retaining other concepts. In contrast, our method achieves 75.3% PSR-1 and 98.0% PSR-5, which is a significant improvement compared to the best baseline, UCE, with 62.1% PSR-1 and 96.0% PSR-5. This result underscores the effectiveness of our method in simultaneously erasing object-related concepts while preserving other unrelated concepts.
Table 2: Erasing object-related concepts.
Method ESR-1$\uparrow$ ESR-5$\uparrow$ PSR-1$\uparrow$ PSR-5$\uparrow$
SD $22.0\pm 11.6$ $2.4\pm 1.4$ $78.0\pm 11.6$ $97.6\pm 1.4$
ESD $95.5\pm 0.8$ $88.9\pm 1.0$ $41.2\pm 12.9$ $56.1\pm 12.4$
CA $98.4\pm 0.3$ $96.8\pm 6.1$ $44.2\pm 9.7$ $66.5\pm 6.1$
UCE $100\pm 0.0$ $100\pm 0.0$ $62.1\pm 34.6$ $96.0\pm 2.9$
Ours $99.2\pm 0.5$ $97.3\pm 1.9$ $75.3\pm 12.0$ $98.0\pm 0.5$
@@
To gain deeper insights into the behavior of our method, we leverage DAAM (Tang et al., 2022) to visualize the attentive attribution maps that depict the interaction between visual and textual concepts in the generated images. DAAM is an emerging technique that interprets how an input word influences parts of the generated image by analyzing the attention maps in the cross-attention module of the Stable Diffusion model.
We first use DAAM to analyze the original SD model’s behavior in generating images with "Cassette Player" and "English Springer" as input prompts as shown in Figure 2. Each test case comprises eight sub-figures, each of which corresponds to a head in the multihead cross-attention module. As depicted in Figure 2, most heatmaps concentrate on the cassette player and the dog’s body, aligning well with the respective textual prompts. Interestingly, the second head does not focus on the cassette player or the dog’s body but instead on the surrounding background.
We then utilize DAAM to visualize the attribution maps of generated images using the same prompts with our method. We find that on the concept to be retained (i.e., "English Springer"), the heatmaps also focus on the dog’s body except for the second head, mirroring the behavior observed in the original SD model. For the concept to be erased (i.e., "Cassette Player"), the heatmaps exhibit a more dispersed pattern, indicating a lack of specific concentration on any distinct region. This observation suggests that the model, under the erasure effect of our method, diverts attention away from the cassette player concept, providing valuable insights into the underlying mechanism of our method.

### v3

@@ 3.2 Knowledge Hiding and Recovery with Prompt - KPOP
$\underset{\theta^{{}^{\prime}}}{\min}\;\mathbb{E}_{c_{e}\in\mathbf{E}}\left[\underbrace{\left\|\epsilon_{\theta^{{}^{\prime}}}(c_{e})-\epsilon_{\theta}(c_{t})\right\|_{2}^{2}}_{L1}+\lambda\underbrace{\left\|\epsilon_{\theta^{{}^{\prime}}}(c_{e},\mathbf{p}_{c_{e}})-\epsilon_{\theta}(c_{e})\right\|_{2}^{2}}_{L2}\right]$ (1)
Knowledge Hiding/Removal
The goal of this stage is to hide the undesirable concepts $c_{e}\in\mathbf{E}$ from the model $\theta^{{}^{\prime}}$. This can be done by optimizing the OP 1, where minimizing the $L_{1}$ loss enforces the model’s output for undesirable concepts closely resembles that of a target concept $c_{t}$ which is usually a neutral concept like ‘a photo’ or empty description ‘ ’ as in previous works [Gandikota et al., 2023, Bui et al., 2024]. However, minimizing only the $L_{1}$ loss often leads to severe degradation in unrelated concepts. To address this, we simultaneously minimize the $L_{2}$ loss which enables the recovery of hidden concepts using an additional prompt $\mathbf{p}_{c_{e}}$ to guide the recovery process. This guided recovery process reduces the model’s reliance on standard textual inputs, effectively transferring the knowledge of hidden concepts to the prompt $\mathbf{p}_{c_{e}}$.
Knowledge Recovery/Transfer
While the objective of the Knowledge Hiding stage has been previously proposed in [Gandikota et al., 2023, Bui et al., 2024], our main contribution lies in the Knowledge Recovery stage which really makes the hidden knowledge recovery possible rather than being removed as in previous works. Since the fine-tuned model $\theta^{{}^{\prime}}$ has been altered from the original model $\theta$ due to the knowledge hiding process, the goal of this stage is to find the prompt $\mathbf{p}_{c_{e}}$ that can recover the undesirable concept $c_{e}$ that has been removed from the model $\theta^{{}^{\prime}}$. This is formulated as the following optimization problem 2 [Gal et al., 2022].
@@
$\mathbf{p}_{c_{e}}=\underset{\mathbf{p}:\|\mathbf{p}-c_{e}\|_{2}\leq\rho}{\text{argmin}}\;\left\|\epsilon_{\theta^{{}^{\prime}}}(c_{e},\mathbf{p})-\epsilon_{\theta}(c_{e})\right\|_{2}^{2}$ (2)
where $\mathbf{p}_{c_{e}}$ is the learnable prompt, associated with the undesirable concept $c_{e}$, and $\rho$ is the constraint on the prompt to make it not too far from the undesirable concept $c_{e}$. The solution $\mathbf{p}_{c_{e}}$ can be efficiently leanred via gradient descent.
Implication on Backdooring T2I models
Since the sanitized model $\theta^{{}^{\prime}}$ satisfies $\epsilon_{\theta^{{}^{\prime}}}(c_{e},\mathbf{p}_{c_{e}})\approx\epsilon_{\theta}(c_{e})$ while $\epsilon_{\theta^{{}^{\prime}}}(c_{e})\neq\epsilon_{\theta}(c_{e})$, the prompt $\mathbf{p}_{c_{e}}$ effectively acts as a backdoor trigger for the sanitized model $\theta^{{}^{\prime}}$. This means that, even though model $\theta^{{}^{\prime}}$ has been trained to hide the concept $c_{e}$, applying the secret key $\mathbf{p}_{c_{e}}$ can restore its generation capability of the undesirable concept $c_{e}$.
As demonstrated in Section 4.2, the quality of the recovered generated images with the key $\mathbf{p}_{c_{e}}$ is comparable to those produced by the original model $\theta$. Thus, the secret key $\mathbf{p}_{c_{e}}$ serves as a hidden access mechanism, allowing the model developers to flexibly control the model’s generation capability or analyze the model’s vulnerability to future backdoor attacks. We provide a detailed description of our method in Section 6.
@@ 4 Experiments
Table 1: Erasing object-related concepts. Ours denote the results with the setting with the knowledge of to-be-preserved concepts.
Method ESR-1$\uparrow$ ESR-5$\uparrow$ PSR-1$\uparrow$ PSR-5$\uparrow$ RSR-1$\uparrow$ RSR-5$\uparrow$
SD $22.0\pm 11.6$ $2.4\pm 1.4$ $78.0\pm 11.6$ $97.6\pm 1.4$ N/A N/A
ESD $95.5\pm 0.8$ $88.9\pm 1.0$ $41.2\pm 12.9$ $56.1\pm 12.4$ N/A N/A
CA $98.4\pm 0.3$ $96.8\pm 6.1$ $44.2\pm 9.7$ $66.5\pm 6.1$ N/A N/A
UCE $100\pm 0.0$ $100\pm 0.0$ $23.4\pm 3.6$ $49.5\pm 8.0$ N/A N/A
UCE $100\pm 0.0$ $100\pm 0.0$ $62.1\pm 34.6$ $96.0\pm 2.9$ N/A N/A
Ours $99.5\pm 0.3$ $98.0\pm 1.9$ $26.6\pm 5.7$ $47.8\pm 5.0$ $72.0\pm 11.2$ $97.2\pm 2.4$
Ours $99.2\pm 0.5$ $97.3\pm 1.9$ $75.3\pm 12.0$ $98.0\pm 0.5$ $71.5\pm 9.7$ $95.3\pm 3.6$
@@ 4.1 General Settings
Standard evaluation protocol for erasing/unlearning concepts in literature is to evaluate the performance of the sanitized model on erasing tasks and preserving tasks. In this paper, because of the new capability of our method, we propose a new additional metric to evaluate the recovering performance of the sanitized model with the secret key.
More specifically, we conduct three sets of experiments, i.e., erasing object-related concepts, erasing artistic style concepts, and erasing unethical content using Stable Diffusion (SD) version 1.4 as the foundation model. We employ the same setting across all methods, i.e., fine-tuning the model for 1000 steps with a batch size of 1, with the Adam optimizer with a learning rate of $1\mathrm{e}{-5}$. We benchmark our method against four baseline approaches, namely, the original pre-trained SD model, ESD [Gandikota et al., 2023], UCE [Gandikota et al., 2024], and Concept Ablation (CA) [Kumari et al., 2023].
@@ 4.2 Erasing Object-Related Concepts
In this experiment, we assess our method’s capability to remove/hide object-related concepts from the foundation model, such as erasing entire object classes like ‘Cassette Player’. We follow the setting in Gandikota et al. [2023], Bui et al. [2024] and use the Imagenette  [^1] https://github.com/fastai/imagenette, a subset of the ImageNet dataset [Deng et al., 2009], which comprises 10 easily recognizable classes, including ‘Cassette Player’, ‘Chain Saw’, ‘Church’, ‘Gas Pump’, ‘Tench’, ‘Garbage Truck’, ‘English Springer’, ‘Golf Ball’, ‘Parachute’, and ‘French Horn’. We erase a set of 5 classes simultaneously, where we generate 500 images for each class and use the pre-trained ResNet-50 [He et al., 2016] to detect the presence of an object in these images. We employ three metrics: Erasing Success Rate (ESR-k): The percentage of generated images with ‘to-be-erased’ classes where the object is not detected in the top-k predictions. Preserving Success Rate (PSR-k): That with ‘to-be-preserved’ classes where the object is detected in the top-k predictions. Recovering Success Rate (RSR-k): That with ‘to-be-erased’ classes when the secret key is provided.
Removing and Preserving Performance.
@@
We select four distinct sets of five classes from the Imagenette dataset for erasure and present the results in Table 1. First, we note that the average PSR-1 and PSR-5 scores across the four settings of the original SD model stand at 78.0% and 97.6%, respectively. This means that 78.0% of the generated images contain the object-related concepts, which are detected in the top-1 prediction, and this number increases to 97.6% when considering the top-5 predictions. These scores highlight the original SD model’s ability to generate images with the expected object-related concepts.
@@
Regarding erasing performance, all baselines achieve very high ESR-1 and ESR-5 scores, with the lowest being 95.5% and 88.9%, respectively. This demonstrates the effectiveness of these methods in erasing object-related concepts, as only a small proportion of the generated images contained the concepts upon detection. Notably, the UCE method achieves 100% ESR-1 and ESR-5, the highest among the baselines. Our method achieves 99.2% ESR-1 and 97.3% ESR-5, which is much higher than the two baselines ESD and CA, and only slightly lower than the UCE method, which is designed specifically for erasing object-related concepts.
@@
However, despite the high erasing performance, the baselines suffer from a significant drop in preserving performance, with the lowest PSR-1 and PSR-5 scores being 41.2% and 56.1%, respectively. This suggests that the preservation task is more challenging, and the baselines are ineffective in retaining other concepts. In contrast, our method achieves 75.3% PSR-1 and 98.0% PSR-5, which is a significant improvement compared to the best baseline, UCE, with 62.1% PSR-1 and 96.0% PSR-5. Compared to the same setting without the knowledge of to-be-preserved concepts (denoted as UCE and Ours), our method still achieves competitive results, with 3.2% higher PSR-1 but 1.7% lower PSR-5 than UCE. This result underscores the effectiveness of our method in simultaneously erasing object-related concepts while preserving other unrelated concepts.
Figure 1: Qualitative results. (1st column) Original SD model. (2nd column) Sanitized by ESD. (3rd column) Sanitized by our method. (4th column) Recovered by our method using the secret key. Several failure cases in ESD demonstrate incomplete erasure, whereas our approach effectively removes the target concepts while enabling precise recovery with the secret key.
Recovering Performance.
In this section, we highlight a unique property of our method: the ability to recover hidden concepts from the sanitized model using a secret key.
As shown in Table 1, our approach achieves 71.5% RSR-1 and 95.3% RSR-5, closely matching the original SD model’s 78.0% and 97.6% success rates for top-1 and top-5 predictions, respectively. This demonstrates that our method effectively preserves the hidden knowledge while restricting unauthorized access.
For qualitative evaluation, Figure 1 presents a visual comparison of generated images across different models. The first column shows the original images produced by the SD model, while the second and third columns depict sanitized outputs generated by ESD and our method, respectively. The fourth column illustrates the recovered images using our secret key mechanism.
Our method successfully removes the undesirable concepts from the public model while still allowing retrieval when needed. Notably, ESD fails to completely erase certain concepts, leading to inconsistencies in its sanitization process. In contrast, our approach not only ensures effective concept removal but also enables seamless recovery via the secret key, as evidenced by the high-quality regenerated images in Figure 1.
These results confirm that the knowledge of hidden concepts remains intact within the sanitized model but is inaccessible to the public through standard textual prompts, ensuring both security and controlled access. This capability is crucial for scenarios where selective access to sensitive information is required while maintaining the integrity of the generative model.
@@
To gain deeper insights into the behavior of our method, we leverage DAAM [Tang et al., 2022] to visualize the attentive attribution maps that depict the interaction between visual and textual concepts in the generated images. DAAM is an emerging technique that interprets how an input word influences parts of the generated image by analyzing the attention maps in the cross-attention module of the Stable Diffusion model.
We first use DAAM to analyze the original SD model’s behavior in generating images with ‘Cassette Player’ and ‘English Springer’ as input prompts as shown in Figure 2. Each test case comprises eight sub-figures, each of which corresponds to a head in the multihead cross-attention module. As depicted in Figure 2, most heatmaps concentrate on the cassette player and the dog’s body, aligning well with the respective textual prompts. Interestingly, the second head does not focus on the cassette player or the dog’s body but instead on the surrounding background.
We then utilize DAAM to visualize the attribution maps of generated images using the same prompts with our method. We find that on the concept to be retained (i.e., ‘English Springer’), the heatmaps also focus on the dog’s body except for the second head, mirroring the behavior observed in the original SD model. For the concept to be erased (i.e., ‘Cassette Player’), the heatmaps exhibit a more dispersed pattern, indicating a lack of specific concentration on any distinct region. This observation suggests that the model, under the erasure effect of our method, diverts attention away from the cassette player concept, providing valuable insights into the underlying mechanism of our method.